# On Future Power Systems Digital Twins: Towards a Standard Architecture

### v1

@@ III-C Data Engineering Component
The Data Engineering component adds to the data processing capabilities of the DT and enables the DT to collect, clean, integrate, and store data from the different sources previously described (see Sec. III-B). Thereby ensuring that all data are well prepared, accessible, and of high quality for all available modules within the other components. The main modules of this component are Data Collection, Data Formatting, Data Cleaning, Data Integration, and Data Storage. Typically, power system operators maintain databases with detailed infrastructure information. The Data Collection module consists of data systems and interfaces that allow the DT to communicate with these databases and measurement devices to retrieve historical and real-time data about power system assets. The general term Data Interface is used as it refers to the point of interaction or communication between different systems, components, modules, and models. The definition of the proper data interface is application- and use-case-specific, but it can include, for instance, a definition of application programming interfaces (APIs), as well as security requirements (e.g., authentication, encryption). The Data Formatting module is in charge of properly formatting the data (e.g., JSON, XML, CSV, binary, etc.) to support all DT modules. For instance, the common information model (CIM) is a widely used standard in power systems for defining a common format to describe assets like generators, transformers, and substations [39, 40]. Operators adhere to CIM standards for consistent asset naming, ensuring interoperability across systems and modules. The Data Cleaning module checks the veracity of the data and complements, corrects, or discards data where needed, ensuring high data quality. For instance, data cleaning algorithms may detect and remove anomalies in power measurements, such as spikes or missing values, to improve the reliability of analysis and decision-making. Power systems comprise various assets, sensors, and control systems, producing vast amounts of heterogeneous data. The Data Integration module cohesively consolidates data from diverse sources like SCADA systems, smart meters, and weather sensors. This comprehensive integration enables the DT to analyze and optimize power systems holistically. Finally, the Data Storage module enables the large volume of raw and processed data to be stored efficiently for further use in all DT modules. Operators typically employ databases, data warehouses, and data lakes, including cloud-based storage.
@@ III-D Data Analytics Component
The Data Analytics component involves the exploration, interpretation, and visualization of data to derive meaningful insights from the power system operators. In general, data analytics operates across four levels: namely descriptive, diagnostic, predictive, and prescriptive. The first level describes what happened in the past regarding a certain event or phenomena. The second level diagnoses why that happened. Third and fourth level respectively predicts what will happen and optimizes operation based on the prediction. Within the proposed ecosystem architecture, the first two levels of data analytics are covered in the Data Analysis and Data Visualization modules. These modules analyze processed data from the data engineering component and output data from the operation and planning component. For instance, detecting anomalies, identifying system states, evaluating control actions, and visualizing performance indicators are among the capabilities of the underlying models. This enables the DT to describe and diagnose the entire system for its operators. The third level of data analytics is handled by the Data Forecasting and Scenario Generation modules, which predict future system states on different timescales. The former operates from multiple days ahead to intraday time scale, while the latter focuses on months to years ahead time scale. Leveraging historical, meteorological, and socio-economic data, these applications employ statistical, machine learning, and heuristic models for forecasting. The fourth level is covered by the modules in the operation and planning component. It enables predictive maintenance and optimized decision making, enhancing system reliability, efficiency, and resilience [41]. Additionally, predictive and descriptive analytics facilitate trend analysis, load forecasting and risk assessment, empowering operators to proactively manage grid operations and address emerging challenges [42].
@@ III-G System of Systems: Multi-Digital-Twin Models
DSOs are responsible for smaller sections of a more extensive network that interfaces with the national transmission network. At the same time, TSOs are responsible for national networks that may go beyond national borders (e.g., electricity interconnection in Europe). Naturally, DTs can be developed for all this infrastructure at different levels by tailoring the DT ecosystem architecture presented in Fig. 3. In this context, we envision a natural fit to the system of system concept [49]. The concept was adapted to the power systems domain by [34], in which DTs of lower-level infrastructure, such as lines and transformers, are interfaced with DTs developed for higher levels, such as DTs of regional distribution networks. A simple example of this concept is presented in Fig. 4. From the DT ecosystem architecture point of view (see Fig. 3), the interface of DTs at lower and higher levels should occur at the High-Fidelity Simulation Models component by replacing the corresponding lower level model by a tailored interface to its DT. Such a tailored interface must share the appropriate data depending on the related DT modules and the embedded power systems’ life cycle phase (see Sec. IV for a more detailed discussion). In this sense, the coupling of multiple DTs via the proposed ecosystem architecture goes beyond the co-simulation concepts already presented for some DT applications [50], as models are kept intrinsic to each DT. Nevertheless, to achieve this, data standardization, interfaces, and data exchange efforts are fundamental and should simplify and pave the way for DT implementation. A notable example in this regard is the utilization of functional mockup units (FMUs), where simulation models are packaged according to a standardized approach, encapsulating the model’s equations and its corresponding solver [51], enabling simulations using different software platforms. Another example of data standardization and interfaces is the open platform of Vorto [52], enabling the integration of IoT devices in DTs. However, to achieve large-scale power system DT implementation requires mature system engineering design methodologies. These can be delivered by working together with the software engineering and computer science communities.
@@ IV Life Cycle of Power System Digital Twins
The proposed ecosystem architecture in Sec. III reconciles the key features of the proposed DT definition in the power systems domain while highlighting its software components as the backbone of such a model. Nevertheless, this ecosystem architecture does not depict how to embed such DTs into the power system life cycle, thus overlooking the temporal dimensions. Unlike other systems with clear end-of-life or end-of-cycle phases, the life cycle of a power system does not have a definitive endpoint. As a fundamental infrastructure that supports society, the long-term goal of a power system is to ensure, with high reliability and resiliency, the supply of energy to end customers. In this context, the original product life cycle management view for which DTs were initially proposed [7, 8] does not fit with the continuous upgrade and expansion of the power systems’ infrastructure. As a result, existing DT definitions fail to consider the additional advantages that DTs can offer to the long-term planning of the power system.
@@
The relevant life cycle of a power system DT comprises a continuous cycle of control, operation, and short- and long-term planning processes performed by the system operators. These processes range from control commands for frequency regulation on a horizon of seconds to capacity management on a horizon of hours, from maintenance recommendations on a horizon of months to policy-making and investment planning on a horizon of decades. Due to the different time resolutions involved, and considering the DT as an integrated ecosystem, it is not clear (i) how the modules (defined in the Operation & Planning component) should interact, and (ii) how data exchange (managed by the Data Engineering component) between models (available at the High-Fidelity Simulation Models component) should occur.
@@
Fig. 5 attempts to reconcile the time dimension of power system DTs based on consultations with DSOs and aligned with their future needs of integration of multiple models. By highlighting data (and information) flows between modules and models, the proposed ecosystem architecture and its key components are linked to the continuous cycle of processes performed by the system operators. Indeed, as data flows from modules used in control to long-term planning, a data downsampling (or data compression) process must be implemented. For instance, the EMT Models (which runs at the second timescale) could be initialized by making use of the Power Flow Models (which runs at the 5- or 15-minute timescales). Similarly, the Security Analysis module (which runs at the hour timescale to assess operational feasibility) could utilize the Power Flow Models. For infrastructure upgrade and expansion, the Capacity Management module (which runs at the hour timescale) could utilize a lower time resolution assessment. For instance, by making use of the Power Flow Models and Thermal Models, and building on top of modules such as DER Control and Observability Analyzer. Integrating and properly interfacing all these models and modules within the same DT ecosystem will bring added benefits to system operators. For instance, investment decisions regarding infrastructure upgrades can be assessed at the seconds timescale within the same ecosystem with minimal technical effort. With these characteristics, we foresee a shift away from the traditional paradigm of segregated operation and planning tasks, enabling decisions made at any stage within the power system’s operational cycle to be promptly evaluated across different time resolutions.
@@
The flow of data between components, modules, and models poses the risk of facilitating error and uncertainty propagation throughout the entire power system life cycle. Due to the intrinsic complexity of power system DT architectures, these are usually disregarded. Nevertheless, they can largely affect decision-making at different power system’s life cycle phases. In the power systems domain, uncertainty is usually considered from the stochastic behavior of the demand and renewable-based generation, enabling the development of complex stochastic and robust mathematical formulations [53]. Although such uncertainty is relevant, in the context of a power system DT, errors and uncertainties from measurements, data downsampling, and errors and inaccuracy from models must also be considered. In this sense, efforts to reduce the gap between simulation and reality have already been recognized in other domains, such as autonomous driving [54]. Nevertheless, efforts in this direction in power systems are not widely known. Considering uncertainty propagation among models with different time resolutions increases the complexity of power system DT developments, raising several important modeling questions yet to be answered. Certainly, this is a future research direction that requires much attention.
@@
V The Role of Artificial Intelligence & Machine Learning in Future Digital Twins
@@
In the context of AI, ML models have been proven to successfully learn from large data sets, finding hidden patterns and enabling the discovery of new knowledge. As a result, AI and ML are revolutionizing different sectors, from drug production [55] to algorithms development [56], showing even above human performance in complex games [57]. Similar disruptions are expected in the energy sector, with AI and ML models potentially enhancing the performance of several modules that are the backbone for power system operation and planning [58]. Although several of these advancements have been already recognized [59], they have not yet been framed to a specific DT architecture. Below, we present our vision (non-comprehensive due to space limitation) of how state-of-the-art AI and ML models (see also Table II) can enhance some of DT components and modules described in Sec. III. The expected enhancements are divided based on the relevant system operator’s processes.
@@
VI Conclusion

### v2

@@ III-C Data Engineering Component
The Data Engineering component adds to the data processing capabilities of the DT and enables it to collect, clean, integrate, and store data from the different sources previously described (see Sec. III-B). Thereby ensuring that all data are well prepared, accessible, and of high quality for all available modules within the other components. The main modules of this component are Data Collection, Data Formatting, Data Cleaning, Data Integration, and Data Storage. Typically, power system operators maintain databases with detailed infrastructure information, such as the parameters and locations of power system assets, switching states, and network topology. The Data Collection module consists of data systems and interfaces that allow the DT to communicate with these databases and measurement devices to retrieve historical and real-time data on power system assets. A critical aspect is the inclusion of spatio-temporal data, which encompasses time-synchronized field measurements from devices such as PMUs, smart meters, and other measurement devices originating from the WAMSs and SCADA systems. Time-synchronization of spatio-temporal data is vital for producing reliable simulation results. When combined with geographic information system (GIS) data from TSOs’ and DSOs’ databases, the data integration module accurately maps all data to the corresponding power system asset within the DT ecosystem. The general term data interface is used as it refers to the point of interaction or communication between different systems, components, modules, and models. The definition of the proper data interface is application- and use-case-specific, but it can include, for instance, a definition of application programming interfaces (APIs), as well as security requirements (e.g., authentication, encryption). The Data Formatting module is in charge of properly formatting the data (e.g., JSON, XML, CSV, binary, etc.) to support all DT modules. For instance, the common information model (CIM) is a widely used standard in power systems for defining a common format to describe assets like generators, transformers, and substations [40, 41]. Operators adhere to CIM standards for consistent asset naming, ensuring interoperability across systems and modules. The Data Cleaning module checks the veracity of the data and complements, corrects, or discards data where needed, ensuring high data quality. For instance, data cleaning algorithms may detect and remove anomalies in power measurements, such as spikes or missing values, to improve the reliability of analysis and decision-making. Moreover, in instances where sensors are malfunctioning or absent, the module has the capability to identify and estimate these values using historical measurement data and proximate measurements to support simulations. The enhancements that AI and ML can provide in these scenarios are detailed in Section V. Power systems comprise various assets, sensors, and control systems, producing vast amounts of heterogeneous data. The Data Integration module cohesively consolidates data from diverse sources like SCADA systems, smart meters, and weather sensors. This comprehensive integration enables the DT to analyze and optimize power systems holistically. Finally, the Data Storage module enables the large volume of raw and processed data to be stored efficiently for further use in all DT modules. Operators typically employ databases, data warehouses, and data lakes, including cloud-based storage.
@@ III-D Data Analytics Component
The Data Analytics component involves the exploration, interpretation, and visualization of data to derive meaningful insights from the power system operators. In general, data analytics operates across four levels: namely descriptive, diagnostic, predictive, and prescriptive. The first level describes what happened in the past regarding a certain event or phenomenon. The second level diagnoses why this happened. The third and fourth level, respectively, predicts what will happen and optimizes operation based on the prediction. Within the proposed ecosystem architecture, the first two levels of data analytics are covered in the Data Analysis and Data Visualization modules. These modules analyze processed data from the data engineering component and output data from the operation and planning component. For instance, detecting anomalies, identifying system states, evaluating control actions, and visualizing performance indicators are among the capabilities of the underlying models. This enables the DT to describe and diagnose the entire system for its operators. The third level of data analytics is handled by the Data Forecasting and Scenario Generation modules, which predict future system states on different timescales. The former operates from multiple days ahead to intraday time scale, while the latter focuses on months to years ahead time scale. Leveraging historical, meteorological, and socio-economic data, these applications employ statistical, machine learning, and heuristic models for forecasting. The fourth level is covered by the modules in the operation and planning component. It enables predictive maintenance and optimized decision making, enhancing system reliability, efficiency, and resilience [42]. Additionally, predictive and descriptive analytics facilitate trend analysis, load forecasting and risk assessment, empowering operators to proactively manage grid operations and address emerging challenges [43].
@@ III-G System of Systems: Multi-Digital-Twin Models
DSOs are responsible for smaller sections of a more extensive network that interfaces with the national transmission network. At the same time, TSOs are responsible for national networks that may go beyond national borders (e.g., electricity interconnection in Europe). Naturally, DTs can be developed for all this infrastructure at different levels by tailoring the DT ecosystem architecture presented in Fig. 3. In this context, we envision a natural fit to the system of systems concept [50]. The concept was adapted to the power systems domain by [35], in which DTs of lower-level infrastructure, such as lines and transformers, are interfaced with DTs developed for higher levels, such as DTs of regional distribution networks. A simple example of this concept is presented in Fig. 4. From the DT ecosystem architecture point of view (see Fig. 3), the interface of DTs at lower and higher levels should occur at the High-Fidelity Simulation Models component by replacing the corresponding lower level model by a tailored interface to its DT. Such a tailored interface must share the appropriate data depending on the related DT modules and the embedded power systems’ life cycle phase (see Sec. IV for a more detailed discussion). In this sense, the coupling of multiple DTs via the proposed ecosystem architecture goes beyond the co-simulation concepts already presented for some DT applications [51], as models are kept intrinsic to each DT. Nevertheless, to achieve this, data standardization, interfaces, and data exchange efforts are fundamental and should simplify and pave the way for DT implementation. A notable example in this regard is the utilization of functional mockup units (FMUs), where simulation models are packaged according to a standardized approach, encapsulating the model’s equations and its corresponding solver [52], enabling simulations using different software platforms. Another example of data standardization and interfaces is the open platform of Vorto [53], enabling the integration of IoT devices in DTs. However, to achieve large-scale power system DT implementation requires mature system engineering design methodologies. These can be delivered by working together with the software engineering and computer science communities.
@@ IV Life Cycle of Power System Digital Twins
The proposed ecosystem architecture in Sec. III reconciles the key features of the proposed DT definition presented in Fig. 2 in the power systems domain while highlighting its software components as the backbone of such a model. Nevertheless, this ecosystem architecture does not depict how to embed such DTs into the power system life cycle, thus overlooking the temporal dimensions. Unlike other systems with clear end-of-life or end-of-cycle phases, the life cycle of a power system does not have a definitive endpoint. As a fundamental infrastructure that supports society, the long-term goal of a power system is to ensure, with high reliability and resiliency, the supply of energy to end customers. In this context, the original product life cycle management view for which DTs were initially proposed [7, 8] does not fit with the continuous upgrade and expansion of the power systems’ infrastructure. As a result, existing DT definitions fail to consider the additional advantages that DTs can offer to the long-term planning of the power system.
@@
The relevant life cycle of a power system DT comprises a continuous cycle of control, operation, and short- and long-term planning processes performed by the system operators. These processes range from control commands for frequency regulation on a horizon of seconds to capacity management on a horizon of hours, from maintenance recommendations on a horizon of months to policy-making and investment planning on a horizon of decades. Due to the different time resolutions involved, and considering the DT as an integrated ecosystem, it is not clear (i) how the modules (defined in the Operation & Planning component) should interact, and (ii) how data exchange (managed by the Data Engineering component) between models (available at the High-Fidelity Simulation Models component) should occur. We elaborate on these challenges in the following section.
IV-A Life Cycle Aligned With TSOs’ & DSOs’ Processes
Fig. 5 attempts to reconcile the time dimension of power system DTs with the DSOs needs of integrating multiple models used for different operation and planning tasks. By highlighting data (and information) flows between models used for tasks located at different time horizons (e.g., seconds, minutes, years, decades), the envisioned DT architecture and its key components are linked to the continuous cycle of processes performed by the system operators. Nevertheless, integrating data with different time resolutions within these processes and their underlying simulation models poses significant challenges. For example, consider the Dutch household’s demand profile with several time resolutions presented in Fig. 6. Using demand profiles with different time resolutions as input for the DT simulation models can alter their output due to the inherent information loss, especially considering the stochastic behavior loss in lower time resolution data. As a result, important dynamics, such as peak demands, can be overlooked, impacting the power system’s operational resilience, for instance, by underestimating the peak consumption. Therefore, selecting the appropriate time resolution for each simulation model, performing data downsampling or upsampling as needed, and ensuring time synchronization of measurement data and simulation models are essential for accurate results. Nevertheless, this research area is usually overlooked in the power system community.
@@ IV-A Life Cycle Aligned With TSOs’ & DSOs’ Processes
As data flows from modules used in control to long-term planning, as depicted in Fig. 5, it allows for a data downsampling (or data compression) process to be implemented. For instance, the EMT Models (which runs at the second timescale) could be initialized by making use of the Power Flow Models (which runs at the 5- or 15-minute timescales). Similarly, the Security Analysis module (which runs at the hour timescale to assess operational feasibility) could utilize the Power Flow Models. For infrastructure upgrade and expansion, the Capacity Management module (which runs at the hour timescale) could utilize a lower time resolution assessment. For instance, by making use of the Power Flow Models and Thermal Models, and building on top of modules such as DER Control and Observability Analyzer. Integrating and properly interfacing all these models and modules within the same DT ecosystem will bring added benefits to system operators. As a result, investment decisions regarding infrastructure upgrades can be assessed at the second timescale within the same ecosystem with minimal technical effort. With these characteristics, we foresee a shift away from the traditional paradigm of segregated operation and planning tasks, enabling decisions made at any stage within the power system’s operational cycle to be promptly evaluated across different time resolutions [54].
Figure 6: A normalized load profile of a Dutch household, presented for several time resolutions. As can be seen, the inherent stochastic behavior of the load demand is diminished when using a lower time resolution representation. As a result, important dynamics, such as peak demands, can be overlooked. Overlooking important dynamics can impact the power system’s operational resilience, for instance, by underestimating the peak consumption.
IV-B Life Cycle Challenges
@@
The flow of data between components, modules, and models displayed in Fig. 5 poses the risk of facilitating error and uncertainty propagation throughout the entire power system life cycle. Due to the intrinsic complexity of power system DT architectures, these are usually disregarded. Nevertheless, they can largely affect decision-making at different power system’s life cycle phases. In the power systems domain, uncertainty is usually considered from the stochastic behavior of the demand and renewable-based generation, enabling the development of complex stochastic and robust mathematical formulations [55]. Although such uncertainty is relevant, in the context of a power system DT, errors and uncertainties from measurements, data up-/down-sampling, and errors and inaccuracy from models must also be considered. In this sense, efforts to reduce the gap between simulation and reality have already been recognized in other domains, such as autonomous driving [56]. Nevertheless, efforts in this direction in power systems are not widely known. Considering uncertainty propagation among models with different time resolutions increases the complexity of power system DT developments, raising several important modeling questions yet to be answered. Certainly, this is a future research direction that requires much attention.
IV-C Expected Benefits for TSOs & DSOs
We propose that the envisioned DT standard definition and ecosystem architecture offer multifaceted benefits for TSOs and DSOs. The advanced data management backbone ensures robust data engineering and analytics capabilities within a single platform, providing TSOs and DSOs with enhanced insights into power system control and operation. Integrating high-fidelity simulation models allows system operators to effectively manage error and uncertainty propagation, facilitating comprehensive analyses across all control, operation, and planning components. This setup enables the development and validation of new services, which can be seamlessly redeployed at the system operator control center premises. Integrating a dedicated communication component enhances coordination and safeguards platform integrity, crucial for mitigating cyberattack risks when interfacing with external parties. Adopting multi-DT systems supports long-term infrastructure planning for combined power systems. By consolidating all simulation models, data analytics, data engineering, operation and planning components, and Grid as a Service into a single ecosystem, system operators can efficiently test, validate, and implement additional modules with minimal effort. Once the backbone is established, managing and expanding the system can become a straightforward procedure.
@@ IV-C Expected Benefits for TSOs & DSOs
V The Role of Artificial Intelligence & Machine Learning in Future Power System Digital Twins
@@
In the context of AI, ML models have been proven to successfully learn from large data sets, finding hidden patterns and enabling the discovery of new knowledge. As a result, AI and ML are revolutionizing different sectors, from drug production [57] to algorithms development [58], showing even above human performance in complex games [59]. Similar disruptions are expected in the energy sector, with AI and ML models potentially enhancing the performance of several modules that are the backbone for power system operation and planning [60]. Although several of these advancements have been already recognized [61], they have not yet been framed to a specific DT architecture. Below, we present our vision (non-comprehensive due to space limitation) of how state-of-the-art AI and ML models can enhance some of DT components and modules described in Sec. III. The expected enhancements of specific algorithms, methods, and learning paradigms are divided based on the relevant system operator’s processes and are summarized in Table II.
@@
VI Steps to Implementation
A structured approach is necessary to implement the envisioned DT ecosystem architecture effectively. First, establish a collaborative framework by forming a consortium that includes TSOs, DSOs, software engineers, computer scientists, and academic researchers. The consortium’s goal should be to develop mature system engineering design methodologies for DT implementation. Second, standardize workflows and data sharing by creating data protocols and APIs to facilitate communication between DT modules and physical systems. Ensure that all relevant parties consistently use data formats like JSON, XML, and CSV, and that asset information adheres to industry standards such as CIM. Third, the data engineering and analytics components must be developed and implemented. This includes data formatting, cleaning, integration, storage, and visualization capabilities. Utilize advanced AI and ML models to enhance data processing, visualization, anomaly detection, and data imputation. Once the system’s backbone is developed, system operators can begin implementing their operation and planning modules. The next step is validation and testing, where the DT ecosystem is tested in a controlled environment using historical data and simulated scenarios to ensure proper operation.
VII Conclusion